### v1

# FLex&Chill: Improving Local Federated Learning Training with Logit Chilling

Abstract
Federated learning are inherently hampered by data heterogeneity: non-iid distributed training data over local clients. We propose a novel model training approach for federated learning, FLex&\&&Chill, which exploits the Logit Chilling method. Through extensive evaluations, we demonstrate that, in the presence of non-iid data characteristics inherent in federated learning systems, this approach can expedite model convergence and improve inference accuracy. Quantitatively, from our experiments, we observe up to 6$\times$ improvement in the global federated learning model convergence time, and up to 3.37% improvement in inference accuracy.
1 Introduction
Federated learning represents a paradigm enabling the training of effective models within distributed environments without explicitly exposing local data McMahan et al. (2017). This approach provides opportunities for communication and computation efficiency during training by sharing model parameters/gradients learned independently from local leaning processes while suppressing raw data transmissions. Furthermore, since training operations take place in a de-centralized manner, federated learning opens the possibility to parallelize the model training process via the assistance of local client devices. It presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities consisting the overall system Park et al. (2023a).
In contrast to conventional centralized training methods, federated learning faces challenges associated with delayed model convergence and limited performance. A fundamental challenge of federated learning arises from the disparity of training data used at each client device. Federated stochastic optimization encounters difficulties in determining optimal parameters, leading to increased communication overhead between local devices and the central server, as well as the need for additional training on (typically resource limited) local devices Zhao et al. (2018).
To address these challenges, a number of prior research have focused on optimizing the local training operations on federated learning clients or on refining the server-side aggregation process. For instance, FedProx Li et al. (2020) controls the number of iterations for each local device, aiming to train models resilient to challenges posed by non-independent and non-iid data environments. SCAFFOLD Karimireddy et al. (2020) achieves expedited convergence and improved model accuracy McMahan et al. (2017) by introducing a correction term during the model aggregation phase to balance the influence of each client. These operations alleviate the problems posed by the non-iid environment, a common dataset configuration for federated learning clients.
This paper proposes a novel model training approach for federated learning, FLex&\&&Chill (Federated Learning EXploiting logit Chilling), that targets to boost the convergence and model performance via temperature scaling, namely logit chilling, during the local model training operations. Note that a typical neural network generates a probability ($q_{i}$) for each class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq. 1) with the temperature $T$ conventionally set to 1. However, it is not uncommon in other application scenarios to exploit even higher temperature values (i.e., $T\in(1,\infty)$), for example, in knowledge distillation Hinton et al. (2015); Touvron et al. (2021) or neural network calibration Guo et al. (2017).
$\xi(z,i,T)=q_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
When employing a temperature $T$ higher than 1, the probability distribution across classes naturally undergoes smoothing due to the narrowed interval between the logits Guo et al. (2017), making such an approach suitable for applications such as knowledge distillation. Notably, despite altering the probabilities, the order of labels’ probability remains unaffected with respect to changes in $T$. While these works successfully explore the impact of varying temperatures for neural network operations, they have mainly focused on applying high temperatures after the model training process, mostly for the inference operations. Raising $T$ to a higher value reduces the confidence of the training operations. This is especially useful to tame DNNs which are typically overconfident.
On the other hand, in federated learning, where disparate training data is common, increasing the model confidence of each individually trained client model (using a lower $T$) prior to their parameter aggregation can be a more effective approach. Our investigation involves measuring both the gradient flow and the data position shift within the representation space to provide guidelines in this unexplored territory.
Specifically, this work targets to show that utilizing fractional temperature values (i.e., $T\in(0,1)$) during the training process enables effective gradient propagation towards the input layer when applied to federated learning data distribution scenarios. Consequently, we observed that lower temperatures facilitate more efficient data updates within the representation space when dealing with non-iid datasets, a common characteristic of federated learning systems. From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improves the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with with low temperatures. We summarize the contributions of this work in three-fold as the following:
  • To the best of our knowledge, this is the first work that explores the impact of integrating temperature scaling during the training process in a federated learning scenario.
  • We offer empirical evidence that demonstrates the performance enhancements of exploiting logit chilling in federated learning scenarios, by examining the gradient flow and data position shifts within the representation space.
  • We introduce a novel federated learning model training approach, FLex&\&&Chill, showcasing the application of low temperature usage during the training process.
2 Background and Related Work
2.0.1 Temperature scaling
In machine learning, the output probability distribution generated by a model serves diverse purposes. Primarily, classification models offer probabilities for each class, showcasing the model’s confidence in its predictions. This aids in assessing how certain the model is in its predictions (i.e., classification outputs) and allows consideration of the likelihood of each class prior to making a final classification decision. Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions Pearce et al. (2021); Park et al. (2023b).
In neural networks, computing probabilities from logits often involves employing softmax operations, as shown in Eq. 1. Here, the temperature parameter ($T>0$) controls the probability distribution’s concentration towards one class, and is used to manage how the categorical probability is dispersed over the possible classes Jang et al. (2016). Notably, this categorical probability aligns with the concept of entropy in information theory, offering a perspective where the temperature parameter governs the level of disorder or uncertainty within the distribution. Specifically, when $T\in(1,\infty)$ is applied, the probability distribution tends to be smoother due to increased $T$. This arises from the diminished differences between logit values, enhancing the distribution’s entropy. In contrast, with $T\in(0,1)$ within the softmax function, the probability distribution becomes sharper, since lower temperatures amplify the differences between logits, effectively polarizing the distribution and concentrating probabilities closer to the maximum logit value. Consequently, the model becomes more confident in its predictions, as it allocates higher probabilities to fewer classes, thereby reducing uncertainty in the distribution.
Therefore, temperature scaling operates as a means to adjust the probability distribution within a neural network-based system. Previous work in Knowledge Distillation Hinton et al. (2015); Touvron et al. (2021) exploit the temperature value by applying a high temperature $T\in(1,\infty)$ to achieve a smoothed probability distribution. This manipulation aims to magnify small values within the distribution, often representing hidden information encapsulated within the model. Additionally, in generative models Shih et al. (2023); Wang et al. (2020), temperature scaling regulates the randomness inherent in model outputs. This technique allows for the control of variability levels in generated samples, enabling adjustments to generated output diversity. Temperature scaling is also used in in model calibration Guo et al. (2017). Applying temperature adjustments to the model’s output probabilities, typically through a scaling factor $T$, aids in aligning these probabilities more accurately with the true likelihoods or uncertainties present in the data. Likewise, previous work have focused on applying high temperature values for various purposes, mostly for the purpose of conservatively better-understanding the full dataset without bias.
2.0.2 Federated learning
In recent years, federated learning has emerged as a promising paradigm in machine learning, particularly in scenarios where data privacy, security, and the decentralization of model training are core concerns. Unlike traditional centralized training approaches, federated learning enables model training across distributed devices or servers while keeping the data localized; thus, addressing privacy challenges associated with centralized data aggregation. This decentralized learning paradigm involves iterative model training where local updates occur on individual devices, and only aggregated model updates are shared with the central server.
Despite its benefits, federated learning poses several challenges. A body of work have tackled the issue of communication overhead stemming from the exchange of frequent model parameter updates. A number of previously proposed work have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead Park et al. (2023a). Furthermore, federated learning systems are prone to facing issues rising from data heterogeneity caused from the diversity or differences in the characteristics, distributions, or formats of data across federated learning clients. Previous work have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations Yu et al. (2021) and spatio-temporal variations Liu et al. (2020).
Figure 1: Distribution of gradient norm observed at the input layer for correctly and incorrectly inferenced samples with varying temperature values. We observe noticeable improvements in gradient flow to the when training with lower temperatures.
3 Design of FLex&Chill
We initiate our discussions by exploring the impact of logit scaling, also known as temperature scaling, providing empirical evidence on the rationale behind our proposed system. Furthermore, in this section, we introduce our system, referred to as FLex&\&&Chill, a novel local training approach designed for efficiently training a federated learning framework.
3.1 Impact of Training with Lower Temperatures
As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with limited exploration into its application during the training process itself. In this section, we present empirical results on three initial experiments specifically crafted to showcase the implications of employing temperature scaling throughout the training phase operations in different dimensions. Through these investigations, our aim is to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training. We conduct our experiments with the CIFAR10 dataset Krizhevsky et al. (2009) and use a 2-layered CNN architecture McMahan et al. (2017) as the baseline neural network architecture.
3.1.1 Gradient flow patterns
Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present sample distribution plot of how the gradients flow for 50K training samples in Figure 1. Note that here we separate the plots for the correctly inferenced samples (in the foward pass) and the incorrect ones to observe their differences. The propagation of gradients to the input layer serves as an indicative metric for the model’s learning efficacy. Successful gradient propagation to the input layer implies that the model effectively utilizes and comprehends information from diverse features within the input data during the training phase Evci et al. (2022).
As shown in Figure 1, the mode of distribution shifts to larger values with decreasing temperatures for both correct and incorrect predictions. One reason behind this mode shift stems from the gradient function characteristics of softmax. Eq. 2 presents the gradient of cross-entropy for the softmax activation with varying temperature $T$ and hard target, one-hot encoded labels. For brevity, we present the final derivation result and leave details in the Appendix. Since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a gradient boost.
$\frac{\partial\xi}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})$ (2)
Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision for a given sample, the loss term significantly increases as $p_{i}\xrightarrow{}0$ (more details later in this section). The double-folded effect accounts for substantial mode shift with thick and long tails in the gradient distribution for lower training temperatures.
3.1.2 Data position shift in the representation space
To better understand the impact of temperature scaling on the training process, we conducted a study into the shift of a data’s position in the representation space (i.e., the distance to the decision boundary). Drawing inspiration from the concept of ‘travel to decision boundary’ as proposed by Kim et al. Kim et al. (2022), we quantified the distance of a data point’s inference result to the decision boundary by determining the magnitude of epsilon ($\epsilon$) value multiplied by the adversarial perturbation Goodfellow et al. (2014) needed to alter the model’s decision.
For this experiment, we configure an initial neural network and perform inference for 1,000 samples to identify the results’ positions on the representation space. Among the 1,000, we select the ones that were misclassified, and recompute their positions after performing one training operation using that specific data point and compute the difference in positions on the representation space with respect to the decision boundary. By doing so, we can observe how drastically one training operation affects the next inference output with respect to different temperatures used in the training phase.
Figure 2: Distributions depicting differences between distances to the decision boundary before and after model updates for varying training temperatures. Notice that lower temperatures show noticeable shift in estimations’ positions on the representation space, suggesting their aggressiveness in modifying the model even with a small number of training samples.
Figure 2 plots our results in the form of distributions depicting differences between distances to the decision boundary before and after model updates. When higher temperatures ($T={2,4}$) are applied, we can notice that the changes in inference results’ positions are not significant (i.e., plots skewed towards 0). Whereas, with a small $T$, we can notice more pronounced shifts in data positions within the representation space. This suggests that using a small $T$ at the softmax operations can actively steer the model towards (potentially) substantial optimization to better align with the training data.
3.1.3 Impact of temperature on individual samples
Figure 3: Probability distribution for single training sample (class index 4) with different training temperatures. A low $T$ hallucinates the model’s probability distribution, causing the neural network to be trained more aggressively even with a small number of samples.
We now examine the impact of different degrees of temperatures on a single sample prediction. Specifically, we take a look at the probability distribution of prediction for a single sample with varying temperature during training process. In Figure 3 we present the the resulting probability distribution for a single training sample (with ground truth index of 4). As the plots show, a higher temperature, $T$=4 in this case, leads to a widely spread probability distribution, and the neural network tends to less confident in prediction. On the other hand, with lower temperatures the information embedded in the data sample becomes more concentrated to the predicted label. If the prediction is incorrect, this increased prediction on the wrong label results in reduced probability on the correct label and increases the loss. This leads to more aggressive changes despite training from a single sample.
In federated learning, clients will train themselves with a small amount of local data (much smaller than typical centralized training). For the models trained at the clients to hold value at the server during the aggregation process, it is important that the local data characteristics are effectively embedded in its model parameters despite the small local dataset size. The aggressive behavior when applying low training temperatures that we see in Figure 3 and throughout our initial experimental results serve as evidence that the core information from each training element can be effectively extracted and learned at the neural network when low temperatures are applied in the federated learning training process.
3.2 FLex&\&&Chill and Logit Chilling
Based on such empirical observations, we hypothesise that training neural networks with lower temperatures can benefit the model convergence and performance for systems exploiting with non-iid data, which is common for federated learning scenarios. In FLex&\&&Chill we term this as logit chilling, as a we ‘chill down’ the temperature at the softmax operations, thereby altering the logit computation derived from the neural network during the training process. The focus of a FLex&\&&Chill-based federated learning system is to exploit logit chilling to train a client-side model with low temperature values. As we observed through the studies above and as our evaluations will show, this has an effect of effectively exploiting crucial information from the training samples to actively alter the model towards the new knowledge.
The lower temperatures used by Logit Chilling directly influences the local training process on each federated learning client. Notably, the seamless integration of Logit Chilling and FLex&\&&Chill with existing federated learning frameworks is a distinct advantage, allowing for a straightforward implementation within already deployed frameworks. Furthermore, we note that FLex&\&&Chill can be used orthogonality with other efforts that accelerate federated learning training operations such as FedProx and SCAFFOLD.
We formulate the operations of FLex&\&&Chill using Algorithm 1. Here, typical federated learning operations take place at both the server and clients except for the fact that the server determines a target $T$ at each round and local model training operations are performed with $T$.
Algorithm 1 FLex&\&&Chill
Data ($D_{1},D_{2},...,D_{N}$) where $D_{i}$ is the user $i$’s local data.
Model ($w^{1}_{t},w^{2}_{t},...,w^{N}_{t}$) where $w_{t}^{i}$ is the user $i$’s model in $t$-th federated round.
Parameter: Temperature $T\in(0,1)$
1:Server executes:
2:Initialize $w_{0}$
3:for each round $t=1,2,...,R$ do
4:    $S_{t}\xleftarrow{}$ (random set of $m$ clients)
5:    for $k\in S_{t}$ do
6:         $w_{t}^{k}\xleftarrow{}$ ClientUpdate($w_{t},T$)
7:    end for
8:    for $k\in S_{t}$ do
9:         $w_{t+1}\xleftarrow{}$ Aggregation($w_{t}^{k}$)
10:    end for
11:end for
12:ClientUpdate($w_{t}$, $T$):
13:$w_{t}^{k}\xleftarrow{}w_{t}$
14:for data, label $\in D_{k}$ do
15:    $\mathbf{z}=w_{t}^{k}(data)/T$
16:    $l(\mathbf{z})=loss(\mathbf{z};label)$
17:    $w_{t}^{k}\xleftarrow{}w_{t}^{k}-\eta\nabla l(\mathbf{z})$
18:end for
19:Return $w_{t}^{k}$ to server
4 Experiment
We evaluate the performance of FLex&\&&Chill using three datasets: (i) FEMNIST from the LEAF database Caldas et al. (2018b), including the Extended MNIST handwriting data Cohen et al. (2017) , (ii) the CIFAR10 dataset, and (iii) the CIFAR100 dataset Krizhevsky et al. (2009). As metrics, we use the average accuracy and training loss observed at the clients, selected to observe both the performance gain and the effectiveness of FLex&\&&Chill in its training operations. Furthermore, we also present the number of federated learning rounds needed to achieve a pre-defined target accuracy.
(a) FEMNIST-DNN
(b) CIFAR10-CNN
(c) CIFAR100-ResNet18
Figure 4: Average federated learning client-side model accuracy for different temperatures used in training. Overall, lower temperatures show higher average model accuracy for federated learning systems.
4.1 Experiment Setup
Dataset. We detail the datasets used in our work as follows.
  • FEMNIST dataset: Sourced from the LEAF dataset, the FEMNIST dataset comprises of 10 handwritten digits (0-9) and 52 characters (26 lowercase and 26 upper case) images contributed by 712 users, totaling 157,132 samples. We selected a subset of 36 users, each with diverse amounts of local data to account for the non-iid environment. We followed the implementations provided by the original work Caldas et al. (2018a).
  • CIFAR10 and CIFAR100 dataset: We use the baseline CIFAR10 and CIFAR100 datasets Krizhevsky et al. (2009) and manually partition the data over 50 federated learning users, with each user holding 1,000 samples.
Model. For each dataset, we utilized distinct model architectures, taking into account the specific challenges and characteristics inherent to each dataset. For FEMNIST, an MLP model featuring four linear layers was employed. We term this combination as “FEMNIST-DNN” in the rest of our evaluations. In the case of the CIFAR10 and CIFAR100 datasets, experiments were conducted using a two-layer CNN model McMahan et al. (2017) and a ResNet18 He et al. (2016), respectively. These two cases are denoted as “CIFAR10-CNN” and “CIFAR100-ResNet18.”
Experiment Details. Each case underwent local training for 300 federated training rounds, utilizing FedAvg McMahan et al. (2017) as the baseline federated learning framework. FedAvg is a widely used framework for federated learning. In each round, the server selects 10 participating clients, employing Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.001 Ruder (2016). Each client trained its model for 10 local epochs, employing a batch size of 16, and utilized cross-entropy loss as the loss function.
4.2 Average Client-side Accuracy
(a) FEMNIST-DNN
(b) CIFAR10-CNN
(c) CIFAR100-ResNet18
Figure 5: Federated learning client-side model training loss observed for different training temperature values. We can notice that overall a lower training loss is observed when lower temperature values are used for model training.
We report the average client-side model accuracy results obtained for models trained with different temperatures in Figure 4. For all three datasets (and the three respective models used for each test case) we see an increasing accuracy trend observed for lower temperatures. Note that for the FEMNIST-DNN case with $T$=0.05, we noticed that the accuracy converges after 120 rounds; thus, perform an early exit for this case. One interesting aspect we can notice is that the accuracy for the $T$=0.05 case, which is the lowest temperature we test for, does not show the best performance for the CIFAR100-ResNet18 experiment. While the $T$=0.05 case shows the best performance of FEMNIST-DNN and CIFAR10-CNN, for CIFAR100-ResNet18, we see a accuracy worse than the $T$=0.5 case. This suggests that training with the lowest possible temperature value is not always ideal, rather this result highlights the importance of identifying a proper training temperature as we will further discuss in Section 5. A closer zoom into the plots suggest that the best performing configuration can change over the training round and can vary for different datasets, but overall, we can notice that exploiting a temperature value lower than “1” in the training process can help improve the classification accuracy.
The average accuracy for all federated learning clients after 300 training rounds presented in Table 1 confirm our observations, indicating that with $T$=0.05 the CIFAR10 case shows a 3.37% improvement in accuracy compared to the $T$=1 case. Even for other cases, the use of a low temperature value show promising results in terms of accuracy. We emphasize once more that this does not mean that the accuracy will always perform better at the lowest temperature. Rather, our results suggest the need to carefully explore the fractional temperature space in the federated learning training process.
4.3 Training Loss
Next we compare the training loss observed at the clients. In Figure 5 we present the average the loss observed at each federated learning training round. Again, we note that we perform early exit for the FEMNIST-DNN case at 120 rounds when $T$=0.05 due to quick model convergence. We can notice there that the use of lower temperatures in the training phase generally helps quickly lower the training loss during the local training phase for federated learning clients. At the same time, when observing the CIFAR100-ResNet18 plots, we can see that the loss for $T$=0.05 does not converge well. In fact, this serves as a reason behind the lower accuracy reported for CIFAR100-ResNet18.
From the quantitative values in Table 1 where we plot the average loss observed at all federated learning clients after 300 training rounds (120 for the FEMNIST-DNN @ $T$=0.05), we can make similar observations. Especially, for FEMNIST-DNN, we see a six-fold improvement in loss values compared to $T$=1. For the CIFAR100-ResNet18 case, however, we notice that the lowest loss convergence is in fact the $T$=0.25 case rather than 0.05. This observation confirms the need to properly consider and explore fractional temperature values for client-side model training.
4.4 Federated Learning Convergence Analysis
FEMNIST-DNN CIFAR10-CNN CIFAR100-ResNet18 Temperature ($T$) Accuracy (%) Loss Num. of Round Accuracy (%) Loss Num. of Round Accuracy (%) Loss Num. of Round $T=4.0$ 73.60 (-5.06) 0.893 299 (0.44×) 37.58 (-3.76) 1.724 299 (0.84×) 28.05 (-0.20) 2.984 288 (0.69×) $T=2.0$ 77.20 (-1.46) 0.749 205 (0.64×) 40.35 (-0.99) 1.638 270 (0.93×) 27.89 (-0.36) 2.918 288 (0.69×) $T=1.0$ 78.66 (0.00) 0.684 132 (1.00×) 41.34 (0.00) 1.594 251 (1.00×) 28.25 (0.00) 2.918 200 (1×) $T=0.5$ 79.78 (+1.13) 0.654 88 (1.50×) 41.52 (+0.18) 1.551 161 (1.56×) 29.77 (+1.52) 2.875 178 (1.12×) $T=0.25$ 79.50 (+0.85) 0.650 57 (2.32×) 43.41 (+2.07) 1.532 132 (1.90×) 30.63 (+2.38) 2.997 152 (1.32×) $T=0.05$ 80.86 (+2.20) 0.643 22 (6.00×) 44.71 (+3.37) 1.528 106 (2.37×) 27.05 (-2.72) 4.228 239 (0.74×)
Table 1: Average classification accuracy, model training loss, and number of federated learning training rounds needed to achieve the maximum accuracy of the $T$=4 case for each dataset-model combination. Quantitative results suggest the effectiveness of exploiting low temperatures for model training in federated learning both in terms of model accuracy and training efficiency.
Centralized - iid Federated - iid Federated - non-iid $T$ Acc. (%) Loss Acc. (%) Loss Acc. (%) Loss 4.0 58.05 1.152 48.22 1.473 37.58 1.724 2.0 63.27 1.041 51.90 1.403 40.35 1.638 1.0 65.18 1.044 54.66 1.325 41.34 1.594 0.5 64.19 1.030 53.19 1.320 41.52 1.551 0.25 63.88 1.092 54.64 1.265 43.41 1.532 0.05 60.98 1.227 47.60 1.701 44.71 1.528
Table 2: Average inference accuracy and model training loss for CIFAR10-CNN with different configurations. The impact of FLex&\&&Chill is prominent for federated learning scenarios with non-iid datasets, whereas its impact is less when elsewise.
Finally, the last columns for each test case in Table 1 we present the number of federated learning training rounds needed to achieve a preset target accuracy for different temperature values using for model training. This number provides an understanding of how efficient the local training processes are when different temperature values are applied. S
[…]
From the results in Table 1 we can notice that with $T$=0.05 the FEMNIST-DNN case achieves the target accuracy after only 22 training rounds. We see similar trends of low temperatures positively affecting the training efficiency throughout our results. This is from the generally aggressiveness of training behaviors when applying lower temperature training. Given a single training sample, while high temperature-based training conservatively absorbs the available information, lower temperatures are more progressive in making model changes. Given the small datasets used for local training and the need to effectively deliver its knowledge to the server for aggregation, we see FLex&\&&Chill as an effective way to train client-side models in the federated learning process.
4.5 Impact of Dataset Dispersion and Federated Learning on FLex&\&&Chill
We now examine the impact of FLex&\&&Chill on different model training configurations. Specifically, using the CIFAR10-CNN test case, we examine two additional configurations: (i) centralized learning with iid data and (ii) federated learning with iid data. Note that the federated learning network is identical to our previous experiments except for the dataset distribution and was tested for 300 rounds. By comparing these results with the non-iid federated learning results, we target to understand the impact of FLex&\&&Chill on federated learning and dataset disparity in greater detail.
As Table 2 shows, applying FLex&\&&Chill, thus a lower $T$ does not induce a positive impact on the performance of centralized learning scenarios. In fact, for centralized learning, controlling $T$ for the training process does not show a noticeable impact. We see the same phenomena for the federated learning system with iid samples as well, suggesting that the performance enhancements we see in Table 1 is not an effect of the federated learning process itself. We do point out that holding iid datasets in federated learning will show overall higher performance compared to non-iid configurations (regardless from $T$), but is not a practical case to assume when deploying federated learning systems. Rather, as we discussed, the impact of exploiting low training temperatures is beneficial when dealing with non-iid datasets.
Figure 6: Accuracy drop observed before and after server-side model aggregation. The aggregation process of client models trained with lower temperatures better retain the individual knowledge.
4.6 Model stability of low temperature training during federated learning aggregation
As discussed in Section 3, low temperatures can expedite local training even when dealing with a small amount of training data. Nevertheless, the effect of low temperature training on the model generalization and aggregation stability of federated learning has not yet been thoroughly studied. To inspect the stability of federated learning aggregation for models trained with low temperatures, we measured the the performance change of each client-side model before and after the server-side aggregation round.
Figure 6 plots the changes in accuracy observed for both participating and non-participating clients at each federated learning round. Specifically, once the model is aggregated, each client computes its accuracy with both the pre-update and the post-update model and computes their difference. Note that pre-update models for clients participating in the round would have updated their local models (very recently) with their own data. Thus, the pre-model performance will be high. However, since each model is trained aggressively over heterogeneous data, the aggregation process could puddle up this knowledge and the aggregated model may lose information. Ideally, with their local information properly aggregated, the performance drop will be minimal. Non-participating clients on the other hand, would hope to have an improved updated model the federated learning process.
As Figure 6 shows the models trained with low temperatures retain the knowledge better than the medium and hot models for participating clients (solid plots groups at bottom). At the early rounds, the models with high temperature show less performance drop, but the fast learning speed of low temperature model amortizes the performance drop, which results in faster learning of the aggregated model. Furthermore, we can also notice that the aggregation process does not interfere with the performance of non-participating clients (grouped on top), implying that federated learning with low temperature training preserves the generalization power.
5 Discussion
$\bullet$ Identifying the optimal temperature. Our study shows the effectiveness of exploiting lower temperatures as a way to enhance federated learning performance. However, an important relevant question remains unanswered: “How do we determine the appropriate or optimal temperature for different applications?” With varying dataset and models that different federated learning application use, the ideal temperature value will also differ. The scope of this work is on identifying and quantifying the potential benefits of lower temperature-based local training in federated learning and we leave the task of identifying per-purpose optimal temperatures to individual system designers and future research.
$\bullet$ Dynamic per-sample temperature scaling. While our paper focuses on understanding the advantages of maintaining a constant low temperature for the training process, this may not be the ideal approach when targeting maximum performance. In reality, varying temperatures can be applied in the training process with respect to their significance Shin et al. (2022); Zeng et al. (2021). That would mean that the training process can exploit lower temperatures for the samples that is “harder” to learn, while higher temperatures are used for those easier to understand. Alternatively, a more course-grain approach of dynamically adjusting the temperature on a per-federated learning round-basis can be another meaningful approach to adapt towards changing local data characteristics.
$\bullet$ Supporting heterogeneous federated learning. While many federated learning frameworks assume homogeneous devices, recent work have pointed out that this is not always the case Park and Ko (2023). In a federated learning network consisting of devices with heterogeneous computing/network capabilities, the issue of “stragglers”, which are devices that take longer to complete local training, becomes important to address. Leveraging insights gleaned from our empirical observations, applying lower temperatures for potential stranglers can be an interesting way of uniforming the local training duration. This would require a prior analysis on the conditions expected at each device and a global consensus on the temperature configuration. We find this as an interesting topic that requires further investigation.
6 Conclusion
This research introduces a novel training approach for federated learning systems, FLex&\&&Chill, aimed at enhancing convergence and improving model performance by exploiting a temperature scaling method termed logit chilling in federated learning local model training operations. Specifically, our work revolves around the implications of employing fractional temperature values ($10<T<$ 1) during the model training process. Our findings, derived from an extensive set of evaluations using three datasets and three widely applied neural networks for federated learning, demonstrate that exploiting the FLex&\&&Chill approach in the client-side model training process holds the potential to expedite model convergence by inducing more pronounced changes despite a limited number of local samples. Consequently, we show that by doing so, we can also achieve an overall higher inference accuracy. While the impact of lower temperature vary with respect to the neural network architecture and dataset, our evaluations reveal up to 6.00$\times$ improvement in federated learning model convergence speed and 3.37% improvement in inference accuracy. We see this work as a first step into devising a more fine-tuned (and optimal) training mechanism for federated learning systems.